# A New Formulation for Zeroth-Order Optimization of Adversarial EXEmples in Malware Detection

### v1

@@ Appendix B Proofs of Theoretical Results
$\mathbb{E}[F_{h_{k}}(v_{k+1})\,|\,\mathcal{F}_{k}]\leq F_{h_{k}}(v_{k})-\gamma_{k}\|\nabla F_{h_{k}}(v_{k})\|^{2}+\frac{L^{3}d^{2}\sqrt{d}}{\ell}\frac{\gamma_{k}^{2}}{h_{k}}.$
@@
$\gamma_{k}\mathbb{E}[\|\nabla F_{h_{k}}(v_{k})\|^{2}]\leq\mathbb{E}[F_{h_{k}}(v_{k})-F_{h_{k}}(v_{k+1})]+\frac{L^{3}d^{2}\sqrt{d}}{\ell}\frac{\gamma_{k}^{2}}{h_{k}}.$
@@
$\sum\limits_{i=\underline{K}}^{\bar{K}}\gamma_{i}\mathbb{E}[\|\nabla F_{h_{i}}(v_{i})\|^{2}]\leq\mathbb{E}[F_{h_{\underline{K}}}(v_{\underline{K}})-F_{h_{k}}(v_{k+1})]+\frac{L^{3}d^{2}\sqrt{d}}{\ell}\sum\limits_{i=\underline{K}}^{\bar{K}}\frac{\gamma_{i}^{2}}{h_{i}}.$
By the definition of $F_{h_{k}}$, we have $F_{h_{k}}(v)\geq\min F_{z}$ for every $v\in\mathbb{R}^{d}$ and $h_{k}>0$. Thus,
$\sum\limits_{i=\underline{K}}^{\bar{K}}\gamma_{i}\mathbb{E}[\|\nabla F_{h_{i}}(v_{i})\|^{2}]\leq F_{h_{\underline{K}}}(v_{\underline{K}})-\min F_{z}+\frac{L^{3}d^{2}\sqrt{d}}{\ell}\sum\limits_{i=\underline{K}}^{\bar{K}}\frac{\gamma_{i}^{2}}{h_{i}}.$
@@ B.2 Proof of Corollary 1
$\eta_{\underline{K}}^{\bar{K}}\leq\Bigg{(}F_{h_{\underline{K}}}(v_{\underline{K}})-\min F_{z}+\frac{L^{3}d^{2}\sqrt{d}}{\ell}\sum\limits_{i=\underline{K}}^{\bar{K}}\frac{\gamma_{i}^{2}}{h_{i}}\Bigg{)}/\Bigg{(}\sum\limits_{\underline{K}}^{\bar{K}}\gamma_{i}\Bigg{)}.$
Due to the choice of $\gamma_{k}=\gamma$ and $h_{k}=h$, we get

### v2

@@ Appendix B Proofs of Theoretical Results
$\mathbb{E}[F_{h}(v_{k+1})\,|\,\mathcal{F}_{k}]\leq F_{h}(v_{k})-\gamma_{k}\|\nabla F_{h}(v_{k})\|^{2}+\frac{L^{3}d^{2}\sqrt{d}}{\ell}\frac{\gamma_{k}^{2}}{h}.$
@@
$\gamma_{k}\mathbb{E}[\|\nabla F_{h}(v_{k})\|^{2}]\leq\mathbb{E}[F_{h}(v_{k})-F_{h}(v_{k+1})]+\frac{L^{3}d^{2}\sqrt{d}}{\ell}\frac{\gamma_{k}^{2}}{h}.$
@@
$\sum\limits_{i=\underline{K}}^{\bar{K}}\gamma_{i}\mathbb{E}[\|\nabla F_{h}(v_{i})\|^{2}]$
$\leq\mathbb{E}[F_{h_{\underline{K}}}(v_{\underline{K}})-F_{h}(v_{k+1})]+\frac{L^{3}d^{2}\sqrt{d}}{\ell}\sum\limits_{i=\underline{K}}^{\bar{K}}\frac{\gamma_{i}^{2}}{h}.$
By the definition of $F_{h}$, we have $F_{h}(v)\geq\min F_{z}$ for every $v\in\mathbb{R}^{d}$ and $h>0$. Thus,
$\sum\limits_{i=\underline{K}}^{\bar{K}}\gamma_{i}\mathbb{E}[\|\nabla F_{h}(v_{i})\|^{2}]\leq F_{h}(v_{\underline{K}})-\min F_{z}+\frac{L^{3}d^{2}\sqrt{d}}{\ell}\sum\limits_{i=\underline{K}}^{\bar{K}}\frac{\gamma_{i}^{2}}{h}.$
@@ B.2 Proof of Corollary 1
$\eta_{\underline{K}}^{\bar{K}}\leq\Bigg(F_{h_{\underline{K}}}(v_{\underline{K}})-\min F_{z}+\frac{L^{3}d^{2}\sqrt{d}}{\ell}\sum\limits_{i=\underline{K}}^{\bar{K}}\frac{\gamma_{i}^{2}}{h}\Bigg)/\Bigg(\sum\limits_{\underline{K}}^{\bar{K}}\gamma_{i}\Bigg).$
Due to the choice of $\gamma_{k}=\gamma$, we get